OpenStudy (anonymous):

Mr. Gallegos is measuring his garage to see if there is enough room to store a rowboat. Which unit is Mr. Gallegos most likely to use when measuring the length and width of his garage floor? Question 3 options: millimeters centimeters meters

OpenStudy (anonymous):

Heeeeeyyyyyyy

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

@Skielerlucas04

OpenStudy (anonymous):

I know the answer :)

OpenStudy (anonymous):

Oh cool then say it

OpenStudy (anonymous):

meters

OpenStudy (anonymous):

awesome thanks

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

could you help with some more?

OpenStudy (anonymous):

The formula for the perimeter of a rectangle is P = 2l + 2w. The length of a rectangle is 3 times its width. Which expression represents the perimeter of the rectangle? Question 4 options: 2 + w(3w) 2 + w(w + 3) 2(3w) + 2w 2(w + 3) + 2w

OpenStudy (anonymous):

mk

OpenStudy (anonymous):

mk lemme try

OpenStudy (anonymous):

alrightyy

OpenStudy (anonymous):

what school is this from?

OpenStudy (anonymous):

k12 (:

OpenStudy (anonymous):

grade?

OpenStudy (anonymous):

10th

OpenStudy (anonymous):

oki :)

OpenStudy (anonymous):

If you dont know it its fine

OpenStudy (anonymous):

i can get it ill keep trying :)

OpenStudy (anonymous):

Thank you (:

OpenStudy (anonymous):

even though im in 9th grade :o

OpenStudy (anonymous):

I think you'll get it haha (:

OpenStudy (anonymous):

i can give you a different question if you want

OpenStudy (anonymous):

:P so what am i trying to find here?

OpenStudy (anonymous):

No i can do this one its ok

OpenStudy (anonymous):

Ok this question is The formula for the perimeter of a rectangle is P = 2l + 2w. The length of a rectangle is 3 times its width. Which expression represents the perimeter of the rectangle? Question 4 options: 2 + w(3w) 2 + w(w + 3) 2(3w) + 2w 2(w + 3) + 2w

OpenStudy (anonymous):

k im finding perimeter

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

might take a while dont worry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Honestly its ok can you please help me with a different question honestly?

OpenStudy (anonymous):

Evaluate 2r(t + 8) if r = 3 and t = 5. Question 5 options: 33 38 65 78

OpenStudy (anonymous):

oki :P

OpenStudy (anonymous):

just help me answer those other questions (:

OpenStudy (anonymous):

do you know this one?

OpenStudy (anonymous):

lemme try it

OpenStudy (anonymous):

78

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

oh wait no

OpenStudy (anonymous):

dont

OpenStudy (anonymous):

ok i got it

OpenStudy (anonymous):

ok what is it?

OpenStudy (anonymous):

Evaluate 2 x 3 (5 + 8) Step 1 : 5 + 8 = 13 Step 2 : 2 x 3 = 6 So 13 x 6 = 78 answer is 78

OpenStudy (anonymous):

awesome thanks!

OpenStudy (anonymous):

np!!!!

OpenStudy (anonymous):

Sergio is s years old. Which expression shows how old he will be 7 years from now? Question 6 options: 7s 7 – s s + 7 7s + 7

OpenStudy (anonymous):

s + 7

OpenStudy (anonymous):

anything else? :)

OpenStudy (anonymous):

thanks im sorry its late

OpenStudy (anonymous):

im sorry for the late reply

OpenStudy (anonymous):

The cost of a circus ticket is $8. There are y people in a group who want to go to the circus. Which of the following expressions describes the total amount of money the group will need to go to the circus? Question 7 options: (8)(2) 8y 8 – y y + 8

OpenStudy (anonymous):

k

OpenStudy (anonymous):

brb (:

OpenStudy (anonymous):

8y

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Which equation represents the sentence? Six less than 3 times a number is 39. Question 8 options: 3 – 6x = 39 3x – 6 = 39 3(x – 6) = 39 6 – 3x = 39

OpenStudy (anonymous):

@imammalik_806

OpenStudy (anonymous):

How many questions are there?

OpenStudy (anonymous):

3x – 6 = 39

OpenStudy (anonymous):

back sorry!

OpenStudy (anonymous):

hey